How many moles of 0.2 M KMnO4 were necessary to cause the color change to occur with the old H2O2 solution?
1
Expert's answer
2019-06-10T04:48:47-0400

It's important to note that hydrogen peroxide solution is acidic. The following reaction will take place:

2 KMnO4 + 3 H2SO4 + 5H2O2 = 2 MnSO4 + K2SO4 + 5O2 + 8H2O

5 H2O2 + 2MnO4- + 6H+ = 5O2 + 2Mn2+ + 8H2O


If the volume of potassium permanganate solution was 0.1 L (or any other value), then number of moles must be:


"n= M \cdot V = 0.2 \frac {moles}{L} \cdot 0.1 L = 0.02 moles"


As the ratio hydrogen peroxide vs permanganate ion is 5:2, then moles of H2O2:


"moles \space of \space H_2O_2 = \frac {5 \cdot 0.02}{2}= 0.05 \space moles"